OpenStudy (anonymous):

-6x+2y=38 6x-8y=-44 for my answer i got x=1 and y =19 its says its wrong help??

OpenStudy (johnweldon1993):

For this one...I would use the OTHER method I didn't use before...elimination This is where you combine the 2 equations *add them* and eliminate 1 variable -6x+2y=38 6x-8y=-44 So when we add these 2 equations...we get -6x + 6x = 0 *cancels out the 'x' 2y + -8y = -6y 38 + -44 = -6 so we have -6y = -6 --------- -6 -6 y = 1 And now just plug that in for 'y' in one of your equations...lets choose the first one -6x+2y=38 becomes -6x+2(1)=38 -6x + 2 = 38 subtract 2 from both sides -6x = 36 Divide both sides by -6 x = -6

OpenStudy (anonymous):

My mistake, y= 1 x= -6 as above

OpenStudy (johnweldon1993):

Now we can even check this Take the fact that y = 1 and x = -6.....and plug them into both equations -6x+2y=38 6x-8y=-44 Becomes -6(-6)+2(1)=38 36 + 2 = 38 38 = 38 This ones correct now for the 2nd one 6(-6)-8(1)=-44 -36 - 8 = -44 -44 = -44 Correct again...so those are your correct answers

OpenStudy (anonymous):

o gezz thanks darlin again :)

OpenStudy (johnweldon1993):

Lol but again...did it make sense? You can really choose which method, *substitution or elimination* you want to use

OpenStudy (anonymous):

i get the checking one but no the whole how u got there whole thing